# On Exact Space-Depth Trade-Offs in Multi-Controlled Toffoli Decomposition

### v1

@@ IV.3 Proving the Lower Bound on Toffoli Depth using Conditionally Clean Ancilla
From Proposition 1, it is evident that the Toffoli depth remains constant over a range of values for $n$. Since, here we are interested in showing the lower bound, we consider $n$ assumes the highest value in the range, and still show that $\log_{2}n$ is strictly less than the corresponding depth, as follows.
@@
The above theorem shows that using the conditionally clean ancilla, the constant factor $c$ can never be reduced to $1$. In the next section, we demonstrate that the exact Toffoli depth in the Clifford + Toffoli decomposition of an $n$-MCT, is lower bounded by $\lceil\log_{2}n\rceil$, regardless of the technique or the number of ancilla qubits used. This bound is achievable through complete binary tree decomposition of $n$-controlled Toffoli gates.
@@ V Tight $\lceil\log_{2}n\rceil$ Lower Bound on Toffoli Depth
To implement an $n$-MCT, we begin with $n$ leaf nodes, where each pair of control qubits is combined using a Toffoli gate, reducing the number of active qubits in each round. At each stage, either two parent nodes or one parent node along with a leftover node from the previous round are further combined into a new parent node. This process continues iteratively until the final Toffoli gate transfers the accumulated information to the root node (the target qubit).
@@
A complete binary tree structure, having $n$ leaf nodes, has a hight $\log_{2}n$, i.e., when $n$ is not a power of $2$. When $n=2^{k}$, for some $k\in\mathbb{N}$, then it becomes a perfect binary tree, having the depth exactly $\log_{2}n=k$. In both the cases, the number of internal nodes, i.e., ancilla qubit is $n-2$. Additionally, the Toffoli count for both these cases becomes $n-1$.
For example, a $32$-MCT can be implemented using $30$ ancilla qubits, with $33$ Toffoli gates, having a minimum Toffoli depth of $5$. Similarly, all $n$-MCT decomposition, with $17\leq n\leq 32$, can be implemented with a minimum Toffoli depth of $5$. Figure 12 provides a schematic diagram of the $7$-MCT circuit decomposition using $5$ ancilla qubits, and $6$ Toffoli gates, with a Toffoli depth of $\lceil\log_{2}7\rceil=3$. In the diagram, the first four Toffoli gates in the first two columns are implemented simultaneously, having depth 1. The next two Toffoli gates are applied sequentially, having Toffoli depth 1, each.
@@
Towards, the conclusion, let us outline a generalized approach that may provide a theoretical understanding. Given the Algebraic Normal Form (ANF) of a Boolean function, $f:\{0,1\}^{n}\rightarrow\{0,1\}^{m}$, naturally the maximum algebraid degree can be $n$. Thus, the reversible quantum circuit implementing $f$, can be designed with a Toffoli depth of $\lceil\log_{2}n\rceil$, utilizing an exponential number of ancilla qubits and $\mathcal{O}(n)$ Clifford gates. This is because, in the ANF, the maximum degree is $n$, which means that $n$ different inputs are ANDed. Thus, an $n$-MCT is enough. Further, the ANF may contain at most $2^{n}$ terms, and that may require exponential number of ancilla qubits. This is achieved by implementing all the required multi-controlled Toffoli (MCT) gates in parallel to compute the non-linear terms of the ANF. This follows from Theorem 3.
This shows that any $n$-degree Boolean function can be implemented with a Toffoli depth of $\lceil\log_{2}n\rceil$. This idea may also be used while considering the cryptanalytic techniques, when the quantum circuits are actually implemented. In this direction, it is understood that the depth may really be quite less for each module to be implemented. Regarding the anicillae, for practical purposes, the ANF contains poly$(n)$ many terms in it, and in such cases the number of ancillae will also be poly$(n)$, instead of the generic exponential bound.

### v2

@@ IV.3 Proving the Lower Bound on Toffoli Depth using Conditionally Clean Ancilla
From Proposition 1, it is evident that the Toffoli depth remains constant over a range of values for $n$. Since here, we are interested in showing the lower bound, we consider $n$ assumes the highest value in the range and still show that $\log_{2}n$ is strictly less than the corresponding depth, as follows.
@@
The above theorem shows that using the conditionally clean ancilla, the constant factor $c$ can never be reduced to $1$. In the next section, we demonstrate that the exact Toffoli depth in the Clifford + Toffoli decomposition of an $n$-MCT is lower bounded by $\lceil\log_{2}n\rceil$, regardless of the technique or the number of ancilla qubits used. This bound is achievable through complete binary tree decomposition of $n$-controlled Toffoli gates.
@@ V Tight $\lceil\log_{2}n\rceil$ Lower Bound on Toffoli Depth
To implement an $n$-MCT, we begin with $n$ leaf nodes, where each pair of control qubits is combined using a Toffoli gate, reducing the number of active qubits in each round. At each stage, either two parent nodes or one parent node, along with a leftover node from the previous round, are further combined into a new parent node. This process continues iteratively until the final Toffoli gate transfers the accumulated information to the root node (the target qubit).
@@
A complete binary tree structure, having $n$ leaf nodes, has a height $\log_{2}n$, i.e., when $n$ is not a power of $2$. When $n=2^{k}$, for some $k\in\mathbb{N}$, then it becomes a perfect binary tree, having the depth exactly $\log_{2}n=k$. In both cases, the number of internal nodes, i.e., ancilla qubit, is $n-2$. Additionally, the Toffoli count for both these cases becomes $n-1$.
For example, a $32$-MCT can be implemented using $30$ ancilla qubits, with $33$ Toffoli gates, having a minimum Toffoli depth of $5$. Similarly, all $n$-MCT decomposition, with $17\leq n\leq 32$, can be implemented with a minimum Toffoli depth of $5$. Figure 12 provides a schematic diagram of the $7$-MCT circuit decomposition using $5$ ancilla qubits, and $6$ Toffoli gates, with a Toffoli depth of $\lceil\log_{2}7\rceil=3$. In the diagram, the first four Toffoli gates in the first two columns are implemented simultaneously, having depth 1. The next two Toffoli gates are applied sequentially, having a Toffoli depth of 1 each.
@@
Towards, the conclusion, let us outline a generalized approach that may provide a theoretical understanding. Given the Algebraic Normal Form (ANF) of a Boolean function, $f:\{0,1\}^{n}\rightarrow\{0,1\}^{m}$, naturally, the maximum algebraic degree can be $n$. Thus, the reversible quantum circuit implementing $f$ can be designed with a Toffoli depth of $\lceil\log_{2}n\rceil$, utilizing an exponential number of ancilla qubits and $\mathcal{O}(n)$ Clifford gates. This is because, in the ANF, the maximum degree is $n$, which means that $n$ different inputs are ANDed. Thus, an $n$-MCT is enough. Further, the ANF may contain at most $2^{n}$ terms, and that may require an exponential number of ancilla qubits. This is achieved by implementing all the required multi-controlled Toffoli (MCT) gates in parallel to compute the non-linear terms of the ANF. This follows from Theorem 3.
This shows that any $n$-degree Boolean function can be implemented with a Toffoli depth of $\lceil\log_{2}n\rceil$. This idea may also be used while considering the cryptanalytic techniques, when the quantum circuits are actually implemented. In this direction, it is understood that the depth may really be quite less for each module to be implemented. Regarding the ancilla, for practical purposes, the ANF contains poly$(n)$ many terms in it, and in such cases, the number of ancilla will also be poly$(n)$ instead of the generic exponential bound.
@@ VI Conclusion
Acknowledgments: The authors like to thank Dr. Alessandro Luongo (CQT, NUS, and Inveriant Pte. Ltd., Singapore) for his valuable inputs while preparing this paper. Subhamoy Maitra acknowledges the funding support provided by the “Information Security Education and Awareness (ISEA) Project phase - III, Cluster - Cryptography, initiatives of MeitY, Grant No. F.No. L-14017/1/2022-HRD”.